OpenStudy (spring98):

Which is the more reasonable measurement of the distance between the tracks on a CD: 1.6 • 10^–3 mm or 1.6 • 10^3 mm? Justify your answer.

OpenStudy (spring98):

@mathlover2014 can you plz help me? i will medal. :)

OpenStudy (welshfella):

the distance between them is too small to be seen so its going to be a lot less than 1 mm

OpenStudy (welshfella):

are you familiar with negative exponents?

OpenStudy (spring98):

so is it this 1.6 • 10^3

OpenStudy (welshfella):

No 10^3 = 1000 mm

OpenStudy (welshfella):

10^-3 = 1/1000 or 0.001

OpenStudy (spring98):

1/1000right?

OpenStudy (welshfella):

No - its one of the 2 options in your questiion. The numbers I gave were only examples.

OpenStudy (welshfella):

- to help you learn the difference between positive and negative exponents.

OpenStudy (spring98):

ya i know but the negative on will be to small so it will be 1.6 • 10^3 mm

OpenStudy (spring98):

@welshfella am i right?

OpenStudy (welshfella):

NO 1.6 * 10^3 = 1,600 mm which is the same as 1.6 meters

OpenStudy (welshfella):

check my first post

OpenStudy (spring98):

so the answer will be 1.6 • 10^–3 mm

OpenStudy (welshfella):

yes

OpenStudy (spring98):

why?

OpenStudy (welshfella):

well have a look at a CD . Can you see the distance between the tracks/

OpenStudy (spring98):

yes

OpenStudy (spring98):

@welshfella would it because the other one would be to big for the tracks of a CD?